Besides 2 mg/L CH4, the groundwater also contains 8.8 mg/L Fe2+. Calculate how much O, is
consumed (in mmol/L) in the oxidation of CH4 to CO2 and of Fe2+ to Fe3+. In order to perform these
calculation you have to derive the redox half reactions.
1
Expert's answer
2020-12-14T14:47:37-0500

CH4 + 2O = CO2 + 4H

Fe2+ + O= Fe3+.

2 mg x 16 /16 = 2.

8.8 x 16/56=2.5.

2+2.5 /16 = 0.28125 mmol/L